OpenStudy (anonymous):

A spherical balloon is inflated with a gas at the rate of 500 cm3/min. How fast is the radius of the balloon increasing at the instant the radius is 30 cm?

OpenStudy (anonymous):

Here are my possible answer: 5/(36 pi) cm/min 2/(18 pi) cm/min 4/(9 pi) cm/min 7/(24 pi) cm/min

OpenStudy (anonymous):

I would determine the initial volume of the sphere.

OpenStudy (anonymous):

In this case, I believe the formula is (4pi (r)3)3, but Ime not sure

OpenStudy (anonymous):

wiki says its (4/3)*(pi)*r^3. Lets try it out. \[\left(\begin{matrix}4 \\ 3\end{matrix}\right)\Pi r ^{3} \] \[\left(\begin{matrix}4 \\ 3\end{matrix}\right)\Pi 30 ^{3} =3770 cm ^{3}\]

OpenStudy (anonymous):

wut, I am not getting the same results. 1 sec

OpenStudy (anonymous):

actually, I am afraid I dont know! Someone help please!

OpenStudy (anonymous):

But I guess, try to find the relation between the radius's rate of change and the rate of change of the sphere getting bigger. Ill be back